# SpaceByte: Towards Deleting Tokenization from Large Language Modeling

### v1

@@ 1 Introduction
Most language models are trained using tokenization, which partitions text into tokens that typically consist of words or subwords. Tokenization is useful because it significantly decreases the inference and training computational costs of large language models. However, tokenization also imposes several disadvantages, including a performance penalty on text distributions different from what the tokenizer was trained on [1]; increased vulnerability to adversarial attacks [2]; worse character-level modeling performance [3], and additional model complexity. [^1] See also Andrej Karpathy’s tweet twitter.com/karpathy/status/1657949234535211009 and video youtube.com/watch?v=zduSFxRajkE&t=6725s on the disadvantages of tokenization.
Recently, MegaByte [4], MambaByte [5], and more [6, 7] have been proposed as new byte-level autoregressive language models that model bytes instead of tokens. (See [8, 9, 10, 11, 12, 13, 14] for encoder and encoder-decoder byte-level modeling.) To address the longer context size resulting from modeling bytes instead of tokens, MegaByte uses multiscale modeling (which is more challenging for autoregressive models than Transformer encoders [15, 16]), while MambaByte uses Mamba blocks [17] instead of Transformer blocks. But although MegaByte and MambaByte have been shown to perform better than a standard byte-level Transformer, to our knowledge, no byte-level autoregressive large language model architecture has been shown to match the performance of tokenized models when controlling for compute costs.
@@
SpaceByte also utilizes multiscale modeling to improve efficiency by grouping bytes into patches. But unlike MegaByte, which uses a fixed patch size, SpaceByte uses a simple rule to dynamically partition the bytes into patches that are aligned with word and other language boundaries. (A similar technique was also explored by Thawani et al. [6].) Our experiments show that this simple modification is crucial for performance, allowing SpaceByte to outperform other byte-level architectures and roughly match the performance of subword Transformers across a variety of text modalities.
@@ 2 SpaceByte
The most common spacelike byte is the space character. Thus, the global blocks are applied most frequently to predict the first character of a word, which we expect is the hardest character to predict in a given word. With fixed patch size (e.g. as in MegaByte), the global blocks are typically inserted in the middle a word, which we expect is inefficient because predicting the rest of the word could likely be more efficiently accomplished using the local blocks. We define continuation bytes to be spacelike so that languages that do not use spaces between words can still benefit from the global blocks between multi-byte characters (e.g. Chinese characters consists of three bytes in UTF-8).
Although this very simple “spacelike” rule is likely not the optimal rule, we find that it works surprisingly well in practice for English text, LaTeX formatted papers, and code. Nevertheless, a critical future direction is to optimize better rules using data rather than our simple heuristic.
@@ 3 Related Work
Word Boundary Using word boundaries to partition patches in autoregressive language modeling was explored by Thawani et al. [6] (and also Edman et al. [12] for encoder-decoder modeling). However, their experiment methodology differed very significantly from ours. While we study large language model perplexity (via bits-per-byte) in compute-limited settings, Thawani et al. [6] study models trained on small datasets ($\sim 5$MB) over 100 epochs and evaluated using word prediction accuracies. The design of their local blocks also differ from SpaceByte as their local blocks do not attend across word boundaries and they use non-causal attention for the local blocks that preceed the global blocks.
@@ 4.2 More Details
We use fairly standard Pre-LN [36, 20, 37] Transformer [38] blocks with no bias terms. Since MegaByte uses Rotary Position Embedding (RoPE) [39], we also use RoPE for all models (which slightly improves the loss). To prevent loss divergences during training, we use qk-layernorm [40, 41] (which we strongly recommend) for all models; i.e. we add an extra layer-normalization to the query and key vectors in the self-attention layers.
All hyperparameters have been carefully tuned using grid and random searches. See Appendices A and B for more details. [^6] Our training code and data reproduction steps can be found at github.com/kjslag/spacebyte
@@ 5 Results
Figure 3: Pareto frontier of the cross-entropy bits-per-byte3 vs FLOPs-per-byte during inference (details in Appendix A.1) for each model architecture (different colors) trained using $10^{18}$ (thin lines) or $10^{19}$ (thick lines) FLOPs on different datasets (on a log-log scale). Lower and to the left is better. SpaceByte (red) outperforms all other byte-level architectures across the entire Pareto frontier for all datasets. SpaceByte roughly matches the performance of the subword Transformer using SentencePiece tokens, and outperforms the subword Transformer using GPT2 tokens.
@@ 7 Conclusion
For future work, it would be fruitful to investigate if Mamba blocks [17] could further improve SpaceByte’s performance. It would also be desirable to improve upon and generalize our simple global block insertion rule.
SpaceByte uses multiscale modeling where the local model operates on bytes while the global model typically operates on words. Another natural extension of our work is to try recursively applying multiscale modeling at even longer scales, such as the sentence or paragraph level.
@@ A.1 FLOPs
Table 6: Training and inference FLOPs-per-token. We calculate the inference FLOPs-per-token in terms of the numbers of parameters ($m$), shown in Table 5. See Section A.1 for symbol definitions.
@@ B.2 Training
Each model was trained using PyTorch on a single 40GB Nvidia A40 and A100 GPUs with mixed-precision (bfloat16 and float32) training and FlashAttention [46, 47]. SpaceByte-793M+184M took the longest to train, requiring about 10 days on a A100 GPU. [^10] We very roughly estimate that additional preliminary and failed experiments not shown in this work required roughly as many FLOPs as the experiments shown in this work.

### v2

@@ 1 Introduction
Most language models are trained using tokenization, which partitions text into tokens that typically consist of words or subwords. Tokenization is useful because it significantly decreases the inference and training computational costs of large language models. However, tokenization also imposes several disadvantages, including performance penalties for languages that are priortizes less by the tokenizer [1, 2, 3]; increased vulnerability to adversarial attacks [4]; worse character-level modeling performance [5], and additional model complexity. [^1] See also Andrej Karpathy’s tweet twitter.com/karpathy/status/1657949234535211009 and video youtube.com/watch?v=zduSFxRajkE&t=6725s on the disadvantages of tokenization.
Recently, MegaByte [6], MambaByte [7], and more [8, 9, 10, 11, 12] have been proposed as new byte-level autoregressive language models that model bytes instead of tokens. (See [13, 14, 15, 16, 17, 18, 19, 20, 21] for encoder and encoder-decoder byte-level modeling.) To address the longer context size resulting from modeling bytes instead of tokens, MegaByte uses multiscale modeling [22, 23, 24], while MambaByte uses Mamba blocks [25] instead of Transformer blocks. But although MegaByte and MambaByte have been shown to perform better than a standard byte-level Transformer, to our knowledge, no byte-level autoregressive large language model architecture has been shown to match the performance of tokenized models when controlling for compute costs.
@@
SpaceByte also utilizes multiscale modeling to improve efficiency by grouping bytes into patches. But unlike MegaByte, which uses a fixed patch size, SpaceByte uses a simple rule to dynamically partition the bytes into patches that are aligned with word and other language boundaries. Our compute-controlled experiments show that this simple modification is crucial for performance, allowing SpaceByte to outperform other byte-level architectures and roughly match the performance of subword Transformers across a variety of text modalities.
@@ 2 SpaceByte
The most common spacelike byte is the space character. Thus, the global blocks are applied most frequently to predict the first character of a word, which we expect is the hardest character to predict [27] in a given word. With fixed patch size (e.g. as in MegaByte), the global blocks are typically inserted in the middle a word, which we expect is inefficient because predicting the rest of the word could likely be more efficiently accomplished using the local blocks. We define continuation bytes to be spacelike so that languages that do not use spaces between words can still benefit from the global blocks between multi-byte characters (e.g. Chinese characters consists of three bytes in UTF-8).
Although this very simple “spacelike” rule is likely not the optimal rule, we find that it works surprisingly well in practice for English text, LaTeX formatted papers, and code. Nevertheless, a critical future direction is to optimize [28, 14, 9] better rules using data rather than our simple heuristic.
@@ 3 Related Work
Word Boundary Prior works have shown utility in using word boundaries to partition patches for autoregressive multi-scale byte-level modeling [9, 8, 11] (and also [18] for encoder-decoder modeling). However, these works did not compare autoregressive byte-level models to subword-level models, nor did they identity a patch partitioning rule that could generically be applied to UTF-8 encoded text. Our primary contributions beyond these prior works is to show how to scale word-boundary byte-level modeling to more diverse text modalities while roughly matching the performance of subword-level models in compute-controlled experiments.
Nawrot et al. [9] and Fleshman and Van Durme [11] make use of the Hourglass Transformer architecture [23]. The SpaceByte architecture is similar to the Hourglass Transformer, except SpaceByte uses a simpler technique for shortening and upscaling the activations before and after the global blocks, and SpaceByte uses a sliding window attention [29, 30, 31] in the local blocks to improve performance for long context sizes.
@@ 4.2 More Details
We use fairly standard Pre-LN [46, 30, 47] Transformer [48] blocks with no bias terms. Since MegaByte uses Rotary Position Embedding (RoPE) [49], we also use RoPE for all models (which slightly improves the loss). To prevent loss divergences during training, we use qk-layernorm [50, 51, 52] (which we strongly recommend) for all models; i.e. we add an extra layer-normalization to the query and key vectors in the self-attention layers.
All hyperparameters have been carefully tuned using grid and random searches. See Appendices A and B for more details. [^6] Our training code and data reproduction steps can be found at github.com/kjslag/spacebyte.
@@ 5 Results
Figure 3: Pareto frontier of the cross-entropy bits-per-byte3 vs FLOPs-per-byte during inference (details in Appendix A.1) for each model architecture trained using $10^{18}$ (connected by thin lines) or $10^{19}$ (thick lines) FLOPs on different datasets (on a log-log scale). Lower and to the left is better. SpaceByte (red) outperforms all other byte-level architectures across the entire Pareto frontier for all datasets. SpaceByte roughly matches the performance of the subword Transformer using SentencePiece tokens, and outperforms the subword Transformer using GPT2 tokens.
@@ 7 Conclusion
Limitations SpaceByte uses a simple byte partitioning rule that relies on “spacelike” bytes, such as spaces which typically denote word boundaries. As such, SpaceByte should not be expected to perform well on arbitrary sequences of bytes, such as images or audio. Some languages, such as Chinese, do not use spaces between words. SpaceByte is somewhat robust to these languages, since e.g. Chinese characters are encoded using three bytes in UTF-8, which SpaceByte will group together. However, our preliminary experiments suggest that SpaceByte performs worse than subword transformers on Chinese text. It would therefore be desirable to improve upon and generalize SpaceByte’s global block insertion rule.
Future Work SpaceByte uses multiscale modeling where the local model operates on bytes while the global model typically operates on words. Another natural extension of our work is to try recursively applying multiscale modeling at even longer scales, such as the sentence or paragraph level. It would also be fruitful to investigate if Mamba blocks [25] could further improve SpaceByte’s performance.
@@ A.1 FLOPs
Table 6: Inference FLOPs-per-token. We calculate the inference FLOPs-per-token in terms of the numbers of parameters ($m$), shown in Table 5. See Section A.1 for symbol definitions.
@@ B.2 Training
Each model was trained using PyTorch on a single 40GB Nvidia A40 and A100 GPUs with mixed-precision (bfloat16 and float32) training and FlashAttention [57, 58]. SpaceByte-793M+184M took the longest to train, requiring about 10 days on an A100 GPU. [^10] We very roughly estimate that additional preliminary and failed experiments not shown in this work required roughly as many FLOPs as the experiments shown in this work.